OpenStudy (anonymous):

Jason ran for 1.5 hours with a 2 mph wind at his back. It took him 2.5 hours to get home with the same wind in his face. If he ran in the same manner and effort, how fast would Jason have run without the wind?

OpenStudy (anonymous):

(1.5 hours)(x+2 mph) = d/2 (2.5 hours)(x-2 mph) = d/2 x = 8 mph d/2 = 15 miles distance/speed = time d =30, speed = 8mph he would have ran for 3.75 hours with no wind

OpenStudy (anonymous):

Thank you! But that's not an option :/ a. 5 mph b. 10 mph c. 15 mph d. 20 mph those are my options.

OpenStudy (anonymous):

b